OpenStudy (anonymous):

if f(x)=2x+1 and g(x)=x+2 then f(g(x)), which is f(g(x))

OpenStudy (aakashsudhakar):

f(x) = 2x + 1; g(x) = x + 2 f(g(x)) = f(x + 2) = 2(x + 2) + 1 = 2x + 4 + 1 = 2x + 5 f(g(x)) = 2x + 5 g(f(x)) = g(2x + 1) = (2x + 1) + 2 = 2x + 3 g(f(x)) = 2x + 3

OpenStudy (anonymous):

f(g(x)) means find the value of f when x has a value of g(x) so f(x+2) = 2(x+2) + 1 = 2x + 5